# VersatileFFN: Achieving Parameter Efficiency in LLMs via Adaptive Wide-and-Deep Reuse

### v1

@@ 4.2 Evaluation Results
To rigorously evaluate the efficacy of our proposed method, we conduct continued pre-training on a pre-trained base model for one additional epoch, utilizing the identical corpus. We benchmark our approach against the following methods: 1) Mixture-of-Experts (MoE): A sparse architecture that inherits the configuration of the dense base model but incorporates additional 8 small experts with activating top-2 experts during the forward pass. 2) $k$-Loop: A variant of the dense architecture that retains the original architecture but introduces recurrence by iteratively applying the FFN $k$ times at each layer.
@@
Efficiency Comparison. We first provide an analysis of the efficiency in Table 2, including the average accuracy, model parameters, and FLOPs of the FFN part. We calculate the FLOPs as 1 token input and 1 token output. Since the same architecture is maintained, the number of parameters for methods $k$-Loop is consistent with that of the base model. The proposed VersatileFFN method, however, introduces a router and loop predictor, which results in a slightly higher parameters. The MoE method, on the other hand, introduces real small experts on top of the base model, leading to a significant increase in parameters. For FFN FLOPs, the computational cost of $k$-Loop method is clearly $k$ times that of base. We calculate the FLOPs of VersatileFFN based on the inference statistics of the ARC-c dataset. Specifically, we calculate the average loops of all layers $N_{mean}$ and the proportion $P$ of loops not equal to the maximum loop, and then compute the FLOPs by $Base\times N_{mean}+(MoE-Base)\times P$.
@@
Benchmark Accuracy. We report the detailed zero-shot performance of VersatileFFN against other methods in Table 1. From the results, VersatileFFN consistently achieves the highest average accuracy across both model scales. Specifically, on the 720M scale, our method attains an average accuracy of 57.03%, outperforming the strong MoE baseline (55.87%) and the 6-Loop method (56.55%). It is worth noting that while MoE achieves the lowest loss (2.411 vs. 2.430 for VersatileFFN), this advantage does not translate into downstream task accuracy. VersatileFFN demonstrates more robust generalization capabilities, particularly on reasoning-intensive tasks such as ARC-e and COMM, where it leads by a significant margin (+3.33% over MoE on ARC-e).
@@
As shown in Table 2, a key advantage of VersatileFFN is its parameter efficiency. Unlike MoE, which significantly increase the total parameters (from 720M to 1145M), VersatileFFN introduces negligible parameter overhead ($0.1\%$ increase). Regarding computational complexity, VersatileFFN is significantly more efficient than the 4-Loop and 6-Loop methods. For instance, at the 350M scale, VersatileFFN requires approximately 45% fewer FLOPs than the 6-Loop method while achieving superior accuracy. This indicates that VersatileFFN allocates computational resources more effectively than simply stacking recurrent loops, offering a compelling balance between model size, training cost, and downstream performance.
@@ 4.3 Visualized Analysis
Actual loops. We first visualize the actual loops executed at each layer on the ARC-c dataset, as shown in Figure 4. We observe distinct behavioral patterns depending on model size: the smaller 354M model tends to allocate a higher computational budget to the final layers (with a sharp increase in layers 11–14), whereas the larger 720M model exhibits a ”middle-heavy” distribution, concentrating recurrence primarily within the intermediate layers.
Figure 4: The average predicted loops per layer on ARC-c dataset. Left: 354M base model. Right: 720M base model.
Features of two branches. We then analyze the features of the width-versatile and depth-versatile at Layer_0 using a random sample from the ARC-c, as visualized in Figure 5. Although both branches inherit the same base parameters, the output features of the two branches are not exact replicas. This demonstrates that they learn complementary representations within a globally aligned semantic space.
Figure 5: Heatmap of cosine similarity at Layer 0. Left: 354M base model. Right: 720M base model.
Word cloud. In Figure 6, we analyze the word cloud of 354M model at Layer_0 on ARC-c. A lower gating coefficients $\lambda$ indicates more loops. The words in the left is dominated by specific action verbs and similar words such as clean, remove, cut, cup. Conversely, the right consists primarily of high-frequency, generic terms like make, use, water, will. These tokens typically rely on shallow surface statistics or syntactic cues, allowing the model to capture their representations with minimal recurrence.
@@ 4.4 Ablation Studies
We conduct a series of ablation studies using the 354M base mdoel on 40B tokens.
Analysis on Each Components. From Table 3, we observe that both width-versatile and depth-versatile individually outperform the baseline. The combination of all modules yields the lowest loss and the highest average accuracy. This result confirms that the two branches are complementary, functioning synergistically to enhance model expressivity without redundancy.
@@
VersatileFFN 2.633 50.98
VersatileFFN 2.625 51.86
VersatileFFN 2.618 52.10
VersatileFFN 2.617 52.33
@@
Table 4: Analysis of the effect of: (A) Various Loops under MoE-8-4; (B) Various experts under 4-Loop.
@@
Performance without Continued Pre-training. Table 5 shows the performance without continued pre-training. VersatileFFN attains an accuracy of 51.14%, yielding a 3.16% improvement over the Base and outperforming the 4-Loop.

### v2

@@ 4.2 Evaluation Results
To rigorously evaluate the efficacy of our proposed method, we conduct continued pre-training on a pre-trained base model for one additional epoch, utilizing the identical corpus. We benchmark our approach against the following methods: 1) Mixture-of-Experts (MoE): A sparse architecture that inherits the configuration of the dense base model but incorporates additional 8 small experts with activating top-2 experts during the forward pass. 2) $k$-Loop: A variant of the dense architecture that retains the original architecture but introduces recurrence by iteratively applying the FFN $k$ times at each layer. To ensure a fair comparison, both methods also share the experimental setting of continued pre-training on the same pre-trained base model.
@@
Efficiency Comparison. We first provide an analysis of the efficiency in Table 2, including the average accuracy, model parameters, and FLOPs of the FFN part. We calculate the FLOPs as 1 token input and 1 token output. Since the same architecture is maintained, the number of parameters for methods $k$-Loop is consistent with that of the base model. The proposed VersatileFFN method, however, designs a router and loop predictor, which introduces negligible parameters. The MoE method, on the other hand, introduces real small experts on top of the base model, leading to a significant increase in parameters. For FFN FLOPs, the computational cost of $k$-Loop method is clearly $k$ times that of base. We calculate the FLOPs of VersatileFFN based on the inference statistics of the ARC-c dataset. Specifically, we calculate the average loops of all layers $N_{mean}$ and the proportion $P$ of loops not equal to the maximum loop, and then compute the FLOPs by $Base\times N_{mean}+(MoE-Base)\times P$.
@@
Benchmark Accuracy. We report the detailed zero-shot performance of VersatileFFN against other methods in Table 1. From the results, VersatileFFN consistently achieves the highest average accuracy across both model scales. Specifically, on the 1.21B scale, our method attains an average accuracy of 60.47%, outperforming the strong MoE baseline (59.65%) and the 6-Loop method (60.05%). It is worth noting that while MoE achieves the lowest loss (2.278 vs. 2.319 for VersatileFFN), this advantage does not translate into accuracy. VersatileFFN demonstrates more robust generalization capabilities, particularly on reasoning-intensive tasks such as ARC-e and COMM, where it leads by a significant margin (+2.81% over MoE on ARC-e).
@@
As shown in Table 2, a key advantage of VersatileFFN is its parameter efficiency. Unlike MoE, which significantly increase the total parameters, for example, from 1.21B to 1.97B (+63%), VersatileFFN introduces negligible parameter overhead. Regarding computational complexity, VersatileFFN is significantly more efficient than the 4-Loop and 6-Loop methods. For instance, at the 350M scale, VersatileFFN requires approximately 45% fewer FLOPs than the 6-Loop method while achieving superior accuracy. This indicates that VersatileFFN allocates computational resources more effectively than simply stacking recurrent loops, offering a compelling balance between model size, training cost, and downstream performance.
@@ 4.3 Visualized Analysis
Actual loops. We first visualize the average predicted loop counts per layer on the ARC-c dataset in Figure 4. The smaller 354M model exhibits a late-stage allocation strategy, dedicating the highest computational budget to the final layers (specifically layers 11–14). The medium-sized 720M model transitions to a middle-heavy distribution, concentrating recurrence primarily within the intermediate layers while tapering off at the boundaries. In contrast, the 1.21B model displays a front-loaded profile: the recurrence intensity peaks sharply at Layer 2 and subsequently stabilizes into a consistent plateau for the remainder of the network.
Figure 4: The average predicted loops per layer on ARC-c dataset. From left to right: 354M, 720M, 1.21B model.
Word cloud. We then analyze the word cloud of 354M model at Layer_0 on ARC-c in Figure 5. A lower gating coefficients $\lambda$ indicates more loops. The words in the left is dominated by specific action verbs and similar words such as clean, remove, cut, cup. Conversely, the right consists primarily of high-frequency, generic terms like make, use, water, will. These tokens typically rely on shallow surface statistics or syntactic cues, allowing the model to capture their representations with minimal recurrence.
@@ 4.4 Ablation Studies
We conduct a series of ablation studies using the 354M base model on 40B tokens.
@@
Ours 2.633 50.98
2.629 51.35
2.618 52.10
2.617 52.33
Analysis on Each Components. From Table 3, we observe that both width-versatile and depth-versatile individually outperform the baseline. When the fusion gating is removed , we adopt simple averaging to fuse the outputs of the width and depth pathways. The combination of all modules yields the lowest loss and the highest average accuracy. This result confirms that the two branches are complementary, functioning synergistically to enhance model expressivity.
@@
Table 4: Analysis of the effect of: (A) Various Loops under MoE-8-2; (B) Various experts under 4-Loop.
@@
Performance without Continued Pre-training. Table 5 shows the performance trained from scratch. VersatileFFN attains an accuracy of 51.14%, yielding a 3.16% improvement over the Base and outperforming other methods.
@@ 5 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Inference Optimization
To minimize latency during inference, we introduce two inference-time optimizations: Discrete Early-Exit and Conditional Parallelism. The complete inference procedure is illustrated in Algorithm 2.
Algorithm 2 Inference Computation of VersatileFFN
0: Output tensor $\mathbf{Y}$
1:Controller & Gating:
2:   Compute post-attention representation $\mathbf{H}$ via Eq. 1
3:   Compute loop probabilities $\mathbf{p}=\text{Softmax}(\mathbf{W}_{loop}\mathbf{H})$
4:   Determine discrete loop count $\hat{\ell}=\arg\max(\mathbf{p})$
5:   Compute difficulty-aware fusion scalar $\lambda$ via Eq. 13
6:Width-Versatile Pathway (Conditional):
7:   if $\lambda>0$ then
8:    Construct virtual experts by slicing shared weights Eq. 5 and Eq. 6
9:    Compute $\mathbf{Y}_{width}$ with Top-$k$ routing via Eq. 7
10:   else
11:    $\mathbf{Y}_{width}\leftarrow 0$ {Prune branch for efficiency}
12:   end if
13:Depth-Versatile Pathway (Discrete Early-Exit):
14:   Initialize $\mathbf{H}^{(0)}\leftarrow\mathbf{H}$
15:   for $\ell=1$ to $\hat{\ell}$
16:    Update recursive hidden state $\mathbf{H}^{(\ell)}$ using full shared weights via Eq. 9
17:   end for
18:   $\mathbf{Y}_{depth}\leftarrow\mathbf{H}^{(\hat{\ell})}$ {Hard selection of final state}
19:Difficulty-Aware Fusion:
20:   Compute output $\mathbf{Y}$ by fusing pathways via Eq. 3
21:Return $\mathbf{Y}$
Appendix B Implementation Details
Experiments Setup. We primarily conduct experiments under a continued pre-training setting. Specifically, we initialize the model using the pre-trained checkpoint and continue training for one epoch on the original corpus. For all models, we set the training sequence length to 4,096 tokens. AdamW optimizer is employed with a weight decay of 0.1 and a gradient clipping threshold of 1.0. The learning rate follows a cosine decay schedule, beginning with a warm-up phase for the first 5% of total steps and subsequently decaying to 10% of the peak value. Regarding model-specific hyperparameters, we use a global batch size of 1.5M tokens with a peak learning rate of $5\times 10^{-4}$ for the 354M model, a global batch size of 1M tokens with a peak learning rate of $4\times 10^{-4}$ for the 720M model and a global batch size of 0.5M tokens with a peak learning rate of $3\times 10^{-4}$ for the 1.21B model. Furthermore, for our method and traditional MoE, we incorporate an auxiliary load balancing loss with a weighting coefficient of $1\times 10^{-5}$.
Evaluation Benchmark Setup. We evaluate our method on a comprehensive set of benchmarks: PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), OpenBookQA (OBQA) (Mihaylov et al., 2018), SciQ (Welbl et al., 2017), ARC-easy (ARC-e) and ARC-challenge (ARC-c) (Clark et al., 2018), CommonsenseQA (COMM)  (Clark et al., 2019), and Winogrande (WINO) (Sakaguchi et al., 2019). We utilize OLMES (OLMo et al., 2024) to assess the performance under zero-shot setting.
Appendix C Features of two branches
We analyze the features of the width-versatile and depth-versatile at Layer_0 using a random sample from the ARC-c, as visualized in Figure 6. Although both branches inherit the same base parameters, the output features of the two branches are not exact replicas. This demonstrates that they learn complementary representations within a globally aligned semantic space.
Figure 6: Heatmap of cosine similarity at Layer 0. From left to right: 354M, 720M, 1.21B model.
Appendix D Evolution of Downstream Accuracy During Training
We illustrate the zero-shot accuracy trajectories across eight benchmarks for three distinct model scales equipped with VersatileFFN in Figure 7, Figure 8 and Figure 9. We observe a consistent positive correlation between compute budget and downstream performance. As the number of training tokens increases, accuracy generally trends upward across all tasks and model sizes. Comparing the three figures reveals clear scaling benefits: the 1.21B parameter model achieves higher absolute accuracy peaks and steeper learning curves compared to its smaller counterparts. Notably, while the global trend is monotonic, the learning curves exhibit high-frequency oscillations, particularly evident in the 720M and 1.21B models on tasks such as ARC-Easy and OpenBookQA.
Figure 7: Accuracy of the 354M model on eight benchmarks as training progresses.
Figure 8: Accuracy of the 720M model on eight benchmarks as training progresses.
Figure 9: Accuracy of the 1.21B model on eight benchmarks as training progresses.